OpenStudy (anonymous):

Find a function F(x) satisfying F ' (x) = sin(6x). F(x)=________ How would I start this? thanks:)

OpenStudy (cwrw238):

integrate sin(6x) with respect to x

OpenStudy (anonymous):

could you please explain how to do that?

OpenStudy (cwrw238):

OK this is an extension of standard form INT sin x dx = - cos x + C INT sin (ax) dx = - cos (ax) * 1/a + C

OpenStudy (cwrw238):

hee r a = 6

OpenStudy (anonymous):

okay, so we plug in? sin(6x)dx= -cos(6x) * 1/6 + C ?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

okay:) what happens now?

OpenStudy (cwrw238):

its usually written (-1/6) cos (6x) + C thats F(x)

OpenStudy (cwrw238):

if you differentiate that it should give sin (6x)

OpenStudy (anonymous):

ohh okay.. so that would be the final answer to my problem? \[F(x)=(\frac{ -1 }{ 6 })\cos(6x) + C \] ? if so, does C just stay like that? or does it stand for something?

OpenStudy (cwrw238):

C is the constant of integration The derivative of a constant is zero so when you do the reverse of differentiation you do not know its value or if there is a constant at all.

OpenStudy (anonymous):

oh so what i wrote above is the final answer? just like that? (did i write it correctly?)

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

oh okay, thank you!!